OpenStudy (anonymous):

the formula of a projectile flying through the air is vsin(theta) and v cos(theta) give me an explanation in how this works

OpenStudy (anonymous):

why does this cause a object to fly through the air

OpenStudy (anonymous):

this is the expression of velocity

OpenStudy (anonymous):

yes i just need a brief explanation in why this works

OpenStudy (anonymous):

if u see any projectile with velocity v launched at an angle t to the horizontal.. u can split its velocity as a vector

OpenStudy (anonymous):

wtdu mean this causes an object to fly?

OpenStudy (anonymous):

well i'm making a calculator program and we have to make an object fly and he gave us this formula and it works so now he wants math behind it idk

OpenStudy (anonymous):

v(x) = vcost v(y) = vsint a(x) = 0, a(y)=g

OpenStudy (anonymous):

whats a(x)=0 and a(y)=g

OpenStudy (anonymous):

a(x) is acceleration along x axis

OpenStudy (anonymous):

a(y) is accn along y-axis which is g= 9.8 m/s^2

OpenStudy (anonymous):

how does this come into play

OpenStudy (anonymous):

x = vcos(theta) . t y = vsin(theta).t - 0.5 gt^2

OpenStudy (anonymous):

u know wht a parabola is?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

now t is the parameter here, eliminate it

OpenStudy (anonymous):

so divide t and get x/t=v cos(theta) and y/t=v sin(theta)-.5gt^2

OpenStudy (anonymous):

find the val of t frm the first eqn and put it into the 2nd

OpenStudy (anonymous):

ull get a parabola

OpenStudy (anonymous):

\[y = xtan{\theta} - gx^{2}/2u^{2}\cos^{2}{\theta}\]

OpenStudy (anonymous):

isnt this a parabola?

OpenStudy (anonymous):

is it or is it not?

OpenStudy (mathteacher1729):

Helpful illustration: http://en.wikipedia.org/wiki/File:Dot_Product.svg Sin of an angle gives you the vertical component of the vector. Cos of an angle gives you the horiontal component of the vector. The notation |A| means "the length of the vector A".